# A Lean Dataset for International Math Olympiad: Small Steps towards Writing Math Proofs for Hard Problems

### v1

@@ 4 Case study: Insights about the dataset
One way to address the evaluation of generalization abilities of an AI model is to test them on challenging problems where the correct answer is not publicly available. For example, FunSearch [23] works on combinatorial problems such as bin packing and comes up with a heuristic algorithm that is better than the best algorithm available in the literature. The last improvement on this problem was made decades ago, so it is clear that this algorithm was not copied from somewhere else.
@@
AlphaGeometry [29], on the other hand, evaluates its performance on all IMO problems in geometry from 2000 to 2022. To avoid training/testing set contamination, it trains its language model on a synthetic dataset generated automatically using an SMT solver. This method of using synthetic data addresses many of the evaluation concerns, but the starting point for creating the synthetic data is geometric figures, and then they exhaustively apply all the possible geometric lemmas to those figures. This approach works for geometric problems, especially when a capable SMT solver is available. Moreover, these geometric problems are decidable unlike the ones we study in this paper. For problems in number theory and algebra, the approach of exhaustively applying all the possible lemmas to a given state is not practical, as the number of applicable lemmas is much larger compared to geometric problems. Moreover, AlphaGeometry considers proving the problems in a system like Lean to be difficult and uses a more simple symbolic system developed by the mathematicians.
@@
LeanDojo [32] observes that: "the common practice of splitting theorems randomly into training/testing has led to an overestimated performance in the previous papers. LLMs can prove seemingly difficult theorems simply by memorizing the proofs of similar theorems during training." To remedy this issue, it proposes a new data set in which the testing set is designed to be challenging compared to the training set.
@@
Outside the domain of formal mathematics, Skill-Mix [33] proposes an evaluation technique in which prompts are designed to include a combination of skills (such as metaphor, red herring, and common knowledge physics). Combining skills in this way aims to prompt the model to come up with a sensible answer that most likely does not already exist in the training set of the model. If the model were to produce a sensible answer, it would need to draw from various parts of its training set. Skill-Mix’s study provides positive evidence that GPT-4 can occasionally provide sensible answers to prompts combining a small number of those skills. But for most prompts, GPT-4 does not succeed.
@@
Other studies use pretrained language models to prove theorems, i.e., LLMs that are trained on the contents of the Web. Some of these papers acknowledge that the answers to some of the testing problems might have leaked into the model training set, for example, [11] reports: "there is the potential for proofs from the test set to have leaked into the LLM pretraining data. While the pretraining data for the Minerva LLM at the base of our models does not include the PISA dataset, it does contain code that may include some Isabelle/HOL proofs found in PISA. This should be kept in mind when interpreting the results."
@@
Some other studies do not try to make a clear separation between the training and testing set; neither do they try to distinguish between the generalization and possible retrievals from the the training set or the prompt. For example, LEGO-prover [30] uses ChatGPT to convert a human written proof into a proof sketch for each of the problems in miniF2F, and then again uses ChatGPT to write the formal proof for each of the steps in the proof sketch. When prompting ChatGPT, it tags the possible relevant lemmas and their proofs in the prompt so that GPT can use them if needed. These additional lemmas are taken from a separate library. For each problem in miniF2F, LEGO-prover performs up to 100 proof attempts with ChatGPT costing around 300 dollars. At the end of the day, it is not clear how many of those success rates are randomly exhausting all possible combination of lemmas and how many of them might be direct retrieval from the contents of the prompts or contents of the ChatGPT’s training set. Moreover, since ChatGPT is a model that evolves over time and it also is not open source, concerns about the reproducibility of the results arise.
@@ 6 Conclusion and Future Work
In this work, we presented a dataset of more than 900 lemmas as a stepping stone towards writing automated formal proofs for challenging IMO problems in number theory and algebra. We devised a systematic method to create a large data set of approachable lemmas formalized in Lean. These lemmas are the building blocks of proofs for IMO problems. They are considerably less challenging than IMO problems, yet GPT-4 struggles to prove almost all of them. Moreover, we performed an extensive human evaluation of GPT-4 responses to diagnose the issues this model encounters when it tries to write a Lean proof for these problems. Our experiments on these lemmas allowed us to diagnose and evaluate the obstacles towards proving the unsolved IMO problems in the miniF2F dataset, and beyond, including the more challenging ones from 2022 and 2023. Developing models that can prove the lemmas in these datasets would brings us closer to models that can participate and do well in IMO.
@@ Appendix A Limitations
Appendix B Other motivations
As the Mathlib library grows, and as more lemmas become available in the library, lemmas that can be utilized in writing new proofs, the amount of time for writing a proof for simple problems is expected to reduce. In writing these proofs, mathematicians use their knowledge of Lean and Mathlib to write the proofs. Machines may also be helpful in reducing the time for the formalization of proofs [6]. GPT-4 has not appeared as a useful tool for writing formal proofs for novel mathematical problems. But, there is a desire and demand for automated systems that are capable to help the mathematicians in writing the formal proofs. In Lean, there are native methods that can suggest tactics or even prove some theorems by applying existing lemmas. Researchers also developed LLMs that can suggest tactics aiming to help mathematicians at writing the formal proofs, e.g., the Copilot by [26]. However, such LLMs appear to be in their infancy, e.g., Copilot only proves easy lemmas that are designed for new learners of Lean.
On the other hand, for mathematicians, there are always harder theorems and conjectures that can be taken up to be formalized. In other words, there is no limit on how large a mathematical library can grow. Most recently, Kevin Buzzard started a project to formalize the proof for Fermat’s last theorem in Lean, and this project is expected to run for a few years [7].
Appendix C Mathematical topics used in our dataset
Table C1 shows the topics used in each of the IMO problems we have studied in this paper. For an automated system to be able to prove the lemmas in our dataset, it will need to utilize these topics and concepts.
Table C1: Topics utilized in the IMO problems formalized and studied in this paper
@@ Appendix C Mathematical topics used in our dataset
Appendix D License and authors’ responsibility statement
Appendix E Decomposing a proof into its building blocks and extracting new lemmas
In this appendix, we explain the procedures and derive some bounds abound extracting new lemmas from an existing proof.
The lemmas in our dataset were created by breaking down the proofs of the IMO problems. These proofs usually have specific structures that we exploit when breaking them down.
Later, we will consider an extreme case and derive some bounds on the number lemmas that can be extracted from an n line proof. But, before that, we explain, how we exploit the structure of the proofs.
E.1 Breaking down the structure of a proof
Let us use the case study in Section 4 as an example. Given a set of hypotheses (explained on page 7), we want to prove $y^{2}$ divides $x$. First, we can consider each of the intermediate hypotheses and define them as a separate lemma. The intermediate hypotheses in the proof are:
  • Lemma 1: prove $y^{2}<x$
  • Lemma 2: prove $2*y^{2}<x$
  • Lemma 4: $(y^{2})^{x}$ divides $x^{x}$
Each of these steps are taken as a lemma since their proofs satisfy the two criteria mentioned before (at least 2 lines of proof and not provable by native provers in Lean).
In an extreme case, a proof may have $k=(n-1)/3$ intermediate hypothesis (each of them with 1 line of statement and 2 lines of proof leading to the 3 in the denominator) which would lead to $2/3*(n-1)$ extracted lemmas. For example, in this setting, a proof with 4 intermediate hypothesis may have $4*3+1=13$ lines. And that leads to $2/3*(13-1)=8$ extracted lemmas, as explained above.
E.2 Case of a proof with no intermediate structure
The forward path would start with the first line of the proof, apply the line, obtain the new proof state, and define that as a new lemma. The proof for this new lemma is the $(n-1)$ lines of the proof that remain. Continuing this way, the forward path can extract $(n-2)$ lemmas where each lemma has at least 2 lines of proof.
E.3 Statistics about our lemmas and their proofs
Table E2: Characteristics about the length of the proof of the lemmas in our dataset
@@ E.3 Statistics about our lemmas and their proofs
Total 907 25,480
Appendix F Datasheet
Following the framework in [12] and similar to [20], Table LABEL:tbl:datasheet provides the data sheet for our dataset.
Table F3: Datasheet for our dataset
@@ Appendix F Datasheet
How many instances are there in total? 907 lemmas.
@@
How was the data associated with each instance acquired? 12 IMO problems were chosen as described in Table 1. The formal proof for each of these problems was written manually by the authors. These proofs were broken into their building blocks leading to a set of 907 lemmas written in Lean 4. The proofs for these lemmas were written by the authors. Lemmas that can be proved by native automated solvers in Lean were excluded from the dataset. The said native solvers are the following: hint, linarith, exact?, simp, omega, ring, norm_cast and norm_num.
What mechanisms or procedures were used to collect the data? Dataset was created by the authors based on the mathematical proofs of 12 IMO problems listed in Table 1.
@@
Was any preprocessing/cleaning/labeling of the data done? No. There was a post processing which excluded lemmas that could be proved directly via native Lean solvers. These lemmas were deemed easy to prove, and since existing solvers can prove them, we do not consider them an obstacle in our larger goal of proving IMO problems.
@@
Is the software that was used to preprocess/clean/label the data available? Lean prover was used to verify the correctness of the process which is an open source system widely used in this field of research.
@@
Has the dataset been used for any tasks already? We have evaluated the ability of GPT-4 in proving the lemmas in our dataset.
Is there a repository that links to any or all papers or systems that use the dataset? There will be a link on GitHub, but we are not sharing it at this time.
@@
Will the dataset be updated? Possibly. If a new version of Lean is released, e.g., Lean 5, we might need to update our dataset based on the new release of Lean and Mathlib.
@@
Appendix G Experimenting with GPT-4: Prompt structure and problem statements
Figure G1 shows an example of how we interact with GPT-4. For each of the lemmas, the prompt starts with the following statement:
Please write a formal proof for this problem in lean language. Please think step by step and explain all your reasoning
After this, the formal lean statement of the lemma is provided.
Figure G1: The structure of the prompt to GPT-4 followed by the formal lemma statement in Lean.
The responses of GPT-4 are graded both for correctness of its explanation in natural language and also for the correctness of its formal proof in Lean. The results were presented in Tables 23.
In cases where GPt-4’s answer is not correct, we follow up with a prompt explaining the mistakes that we detect in the answer, instructing GPT-4 to correct the mistakes. Figures G2 and G3 show examples of such interactions.
Figure G2: Example of hinting GPT-4 and giving feedback about its mistakes.
These interactions rarely lead to a correct answer from GPT-4 as mistakes tend to repeat even after they are explained to GPT-4, e.g., Figure G3. It appears that many of the mistakes and hallucination of lemmas are a byproduct of GPT-4 not having any frame of reference in its training set. Sometimes, after a few rounds of feedback, GPT-4 stops giving any answer and it explains that the problem is hard to prove. Table 4 presents the results of hinting GPT-4 up to 10 rounds.
For these cases, a sensible analogy might be a student who has not studied the materials and shows up at an exam. Such student might improvise and make up answers, answers that might be often nonsensical, drawing from nonexistent rules, principles, and theorems. Such student might also not provide any answers and explain that they are uneducated on the topics of the exam.
Figure G3: Example of hinting GPT-4 and giving feedback about its mistakes.
Appendix H Source Criticism
In the supplementary material, we show the results of using source criticism to compare the correctness of the GPT-4 responses with proofs available on the Internet. We report a strong statistical correlation between the availability of proofs on the Web and the correctness of GPT-4’s responses. To search the Web, we use the Google search engine.
Source criticism is a scientific method and research methodology used in many fields, including information sciences, empirical sciences, and other fields, especially those that work with text [13]. Since the advent of the Internet, a large amount of text has accumulated on the Web. Sometimes it is useful or necessary to trace a given piece of text back to its sources, such as in the case of fact-checking to overcome misinformation, or in legal cases related to copyright. When dealing with the large amounts of text on the Internet, source criticism has proven to be a useful method [27]. We believe this method can also be useful in evaluating the capabilities and generalization of LLMs.
Comparing current texts with previously created texts yields insight into the composition of the newer text. The comparison can tell whether a present text has copied from earlier text and if so how closely the copied text resembles the earlier text. The oldest known variant that informs the later texts qualifies as the primary source. Texts copied or derived from the primary source are secondary sources. A secondary source is not necessarily an exact copy of a primary source. Indeed, a secondary source/text may have novelty of its own by putting together text from two or more different primary sources, or by significantly altering a primary text.
In our case of theorem proving, the primary sources are relatively easy to locate online. For example, most of the publicly available formal proofs written in Lean language can be found on GitHub. LLMs such as ChatGPT [18] and Llama [28], and even specialized models like Llemma [4], ProofNet [3], and Code Llama [24] are trained on contents of the web: either directly trained on Common Crawl or fine tuned from a model that was previously trained on Common Crawl. Although the GPT-4 training set is kept as a secret, we know that previous iterations of this model, such as the GPT-f model, were trained on the Common Crawl and WebMath datasets [21]. The WebMath contains all publicly available code and text on GitHub, arXiv, and Math StackExchange. Hence, the content of the Web can be a reasonable proxy for their training set. Given this, source criticism presents a largely untapped opportunity to evaluate the generalization capabilities of LLMs.
When an LLM produces a proof for a given lemma merely by copying the proof from its training set, we can consider that as a successful direct retrieval. We do not consider such retrieval to be trivial or useless. In fact, correct retrieval of a proof can be a useful capability for a model.
In addition to direct retrieval from the training set, a model may also be able to draw from various sources in its training set and put together a correct response. This would be a more advanced capability compared to direct retrieval and we call that generalization. [^1] Sometimes, generalization is contrasted against memorization, in the machine learning literature and also in educational studies. Direct retrieval that we described here can be considered memorization. This might imply lack of understanding when it comes to learning. However, here, we prefer not to use the term memorization as it may create certain confusions in the context of machine learning, especially the way memorization is used in computer vision. To avoid confusions, we use the term direct retrieval. The source criticism method described above can distinguish between these cases. If a proof (or a piece of text) is a direct copy of a primary source, we can consider that a direct retrieval. However, when a proof for a given lemma does not exactly appear in a primary source, we consider that a more advanced form of behavior by the model and it would no longer be counted as direct retrieval. We consider this form of generalization in three forms:
  1. 1.
    The proof consists of various pieces that are taken from two or more primary sources. The model has taken these pieces from those sources and has managed to put them together in a sensible way.
  2. 2.
    The proof generated by the model does not exactly appear in any primary source, rather, it is an alteration of a proof from a primary source.
  3. 3.
    A combination of the above two forms. For example, a model has retrieved some segment of a proof from some primary source and has put it together with an alternation of some other proof segment.
Here, we only study the correct proofs written by GPT-4, and we would like to know what fraction of these correct proofs can be direct retrieval from its training set (considered to be the Web) and how many of those correct proofs are of the generalization form where the model has made some effort in generating the proof. This effort can be the change of the variable names to align the proof with the current lemma, or it can involve more complex modifications. The results of this analysis is presented in Table H4.
Table H4: Results of analyzing the sources of correct proofs written by GPT-4. An entry of "a/x" in the "direct retrieval" column implies that there were $x$ number of lemmas that GPT-4 were able to write a correct proof for them, and there were $a$ lemmas that can be considered a direct retrieval. Entries of -/0 refer to cases where GPT-4 were not able to write any correct proofs.
# Problem Direct retrieval Modifications from one primary source Complex generalization
1 1959-P1 -/0 -/0 -/0
2 1960-P2 0/1 1/1 0/1
3 1962-P2 1/5 3/5 1/5
4 1964-P2 0/1 1/1 0/1
5 1965-P2 2/8 5/8 1/8
6 1983-P6 -/0 -/0 -/0
7 1984-P6 -/0 -/0 -/0
8 1992-P1 -/0 -/0 -/0
9 1997-P5 -/0 -/0 -/0
10 2022-P2 -/0 -/0 -/0
11 2022-P5 -/0 -/0 -/0
12 2023-P4 -/0 -/0 -/0
Total count 3/15 10/15 2/15
Total percentage 20% 67% 13%
The training set of GPT-4 is not limited to the contents of the Internet. According to its technical report, its training set also contains human generated material, as well as contents obtained from user interactions with GPT-4 accumulated over time. Hence, when the model provides a correct proof and that proof is not found on the Internet, it may still be a direct retrieval from some part of the model’s training set that is not on the Web. This distinction is hard to detect and verify but when interpreting the results in Table H4, it would be useful to keep in mind that the cases of complex generalization might belong to other columns.
One may use source criticism to analyze the incorrect proofs as well. However, given the large amount of incorrect proofs that GPT-4 generated in our experiments, performing source criticism on them may be more labor intensive than insightful. Nevertheless, these are the failure modes of an AI model and studying these failure modes will be crucial when developing automated systems for theorem proving which falls outside the scope of the current paper.
Appendix I Proof repair
For proof repair, we take the incorrect proofs provided by GPT-4 and repair them manually. The goal is to see whether GPT-4’s proof can be useful starting points so that humans do not need to write the proof from scratch. To quantify the effort, we report different measures:
  1. 1.
    Percentage of lines that need to be edited
  2. 2.
    Whether the proof sketch by GPT-4 is correct
  3. 3.
    Levenshtein ratio between GPT-4’s proof and the correct proof
For item 1, if new lines have to be added, those will be counted towards the number of edited lines. As a result, this percentage can become larger than 100%, indicating that much is missing in the generated proof. The second measure aims to identify whether one can take the sketch of the proof from GPT-4 and fill in the structure. The third measure is easy to calculate and from it, we can see a correlation between this measure and the human effort needed to repair the proof (item 1). It can be useful for the community to look at this measure as a proxy for the effort of repairing proofs.
The results will be included in the supplementary material. In summary, for almost all of the lemmas, the effort of writing the proof from scratch is less than the effort needed to go through the GPT-4’s proof and fix the issues in it.

### v2

@@ 4.2 Case of a proof with no intermediate structure
One way to address the evaluation of generalization abilities of an AI model is to test them on challenging problems where the correct answer is not publicly available. For example, FunSearch (Romera-Paredes et al., 2024) works on combinatorial problems such as bin packing and comes up with a heuristic algorithm that is better than the best algorithm available in the literature. The last improvement on this problem was made decades ago, so it is clear that this algorithm was not copied from somewhere else.
@@
AlphaGeometry (Trinh et al., 2024), on the other hand, evaluates its performance on all IMO problems in geometry from 2000 to 2022. To avoid training/testing set contamination, it trains its language model on a synthetic dataset generated automatically using an SMT solver. This method of using synthetic data addresses many of the evaluation concerns, but the starting point for creating the synthetic data is geometric figures, and then they exhaustively apply all the possible geometric lemmas to those figures. This approach works for geometric problems, especially when a capable SMT solver is available. Moreover, these geometric problems are decidable unlike the ones we study in this paper. For problems in number theory and algebra, the approach of exhaustively applying all the possible lemmas to a given state is not practical, as the number of applicable lemmas is much larger compared to geometric problems. Moreover, AlphaGeometry considers proving the problems in a system like Lean to be difficult and uses a more simple symbolic system developed by the mathematicians.
@@
LeanDojo (Yang et al., 2023) observes that: “the common practice of splitting theorems randomly into training/testing has led to an overestimated performance in the previous papers. LLMs can prove seemingly difficult theorems simply by memorizing the proofs of similar theorems during training.” To remedy this issue, it proposes a new data set in which the test set is designed to be challenging compared to the training set.
@@
Outside the domain of formal mathematics, Skill-Mix (Yu et al., 2023) proposes an evaluation technique in which prompts are designed to include a combination of skills (such as metaphor, red herring, and common knowledge physics). Combining skills in this way aims to prompt the model to come up with a sensible answer that most likely does not already exist in the training set of the model. If the model were to produce a sensible answer, it would need to draw from various parts of its training set. Skill-Mix’s study provides positive evidence that GPT-4 can occasionally provide sensible answers to prompts combining a small number of those skills. But for most prompts, GPT-4 does not succeed.
@@
Other studies use pretrained language models to prove theorems, i.e., LLMs that are trained on the contents of the Web. Some of these papers acknowledge that the answers to some of the testing problems might have leaked into the model training set, for example, First et al. (2023) reports: “there is the potential for proofs from the test set to have leaked into the LLM pretraining data. While the pretraining data for the Minerva LLM at the base of our models does not include the PISA dataset, it does contain code that may include some Isabelle/HOL proofs found in PISA. This should be kept in mind when interpreting the results.”
@@
Some studies do not try to make a clear separation between the training and tes sets; neither do they try to distinguish between the generalization and possible retrievals from the the training set or the prompt. For example, LEGO-prover (Wang et al., 2024) uses ChatGPT to convert a human written proof into a proof sketch for each of the problems in miniF2F, and then again uses ChatGPT to write the formal proof for each of the steps in the proof sketch. When prompting ChatGPT, it tags the possible relevant lemmas and their proofs in the prompt so that GPT can use them if needed. These additional lemmas are taken from a separate library. For each problem in miniF2F, LEGO-prover performs up to 100 proof attempts with ChatGPT costing around 300 dollars. At the end of the day, it is not clear how many of those success rates are randomly exhausting all possible combination of lemmas and how many of them might be direct retrieval from the contents of the prompts or contents of the ChatGPT’s training set. Moreover, since ChatGPT is a model that evolves over time and it is not open-source, concerns about the reproducibility of the results arise.
@@ 6 Conclusion and Future Work
In this work, we provided the proofs for all the remaining IMO problems in the test set of miniF2F. We also presented a dataset of 1,329 lemmas as a stepping stone towards writing automated formal proofs for challenging IMO problems in number theory and algebra, derived from a subset of those IMO problems. We devised a systematic method to create a large data set of approachable lemmas formalized in Lean. These lemmas are the building blocks of proofs for IMO problems. They are considerably less challenging than IMO problems, yet state-of-the-art models struggle to prove most of them. Moreover, we performed an extensive human evaluation of o3-mini’s responses to diagnose the issues this model encounters when it tries to write a Lean proof for these problems. Our experiments on these lemmas allowed us to diagnose and evaluate the obstacles towards proving the unsolved IMO problems in the miniF2F dataset, and beyond, including the more challenging ones from 2022 and 2023. Developing models that can prove the lemmas in our dataset would brings us closer to models that can participate and do well in IMO.
Acknowledgments
Xuenan Cao’s work is supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14602223.
@@ Appendix A Limitations
Appendix B Mathematical topics used in our dataset
Table B1 shows the topics used in each of the IMO problems we have studied in this paper. For an automated system to be able to prove the lemmas in our dataset, it will need to utilize these topics and concepts.
Table B1: Topics utilized in the IMO problems formalized and studied in this paper
@@ Appendix B Mathematical topics used in our dataset
8 1985-P6 Natural, Real, NNReal functions, limits, infinite sequences, upper bounds and lower bounds, bijectivity, injectivity, monotonicity, continuity, sets, induction
@@
Appendix C Information about the formal proofs of IMO problems in our dataset
Table C2 lists the problem topic for each IMO problem and whether these proofs appeared in the MiniF2F dataset. We highlight that the majority of the released proofs were not available to the general research community.
Table C2: IMO problems formalized in this paper
# Year Problem Topic in miniF2F Lean proof publicly available # of lines of Lean4 code
1 1959 P1 number theory Yes Yes 9
2 1960 P2 algebra Yes Yes 40
3 1962 P2 algebra Yes No 60
4 1963 P5 algebra Yes No 50
5 1964 P2 algebra Yes Yes 50
6 1965 P2 algebra Yes No 210
7 1968 P5 algebra Yes No 30
8 1969 P2 algebra Yes No 150
9 1974 P3 number theory Yes No 510
10 1981 P6 algebra Yes No 40
11 1982 P1 algebra Yes No 75
12 1983 P6 algebra Yes No 180
13 1984 P6 number theory Yes No 380
14 1985 P6 number theory Yes No 1,310
15 1992 P1 number theory Yes No 480
16 1997 P5 number theory Yes No 390
17 2007 P6 algebra Yes No 570
18 2022 P2 algebra No No 260
19 2022 P5 number theory No No 640
20 2023 P4 number theory No No 450
total 5,884
Appendix D Information about the the extracted lemmas and their formal proofs
Table D3 shows statistics about the length of the proofs for all the lemmas that we have extracted from the IMO proofs, defining our lemma dataset. These lemmas are the building blocks of IMO proofs, and we have used them to evaluate the LLMs.
Table D3: Statistics about the length of the proofs of lemmas in our dataset
@@ Appendix D Information about the the extracted lemmas and their formal proofs
Total 1,329 40,170
Appendix E Additional results on the langauge models
Table E4 presents the accuracy of all the LLMs that we have evaluated on our lemmas dataset. We include additional results on Llama 3.1 (Dubey et al., 2024) and Qwen2.5-Math (Yang et al., 2024) models. We note that these models were not explicitly trained on Lean 4 data but rather perform informal mathematical reasoning. Following this fact, they are unable to generate formal proofs for our lemmas dataset.
Table E4: Comparing performance of SOTA models on lemmas dataset
Problem # of lemmas Deepseek Prover-v1.5-RL (@32) Goedel-Prover (@32) ReProver retrieval ✗ ReProver retrieval ✓ o3-mini (with 10 e.f.) Qwen2.5-Math LLama 3.1
1959-p1 4 3 2 2 2 2 1 0
1960-p2 9 7 6 3 4 1 0 0
1962-p2 14 13 12 7 8 6 0 0
1964-p2 9 5 5 5 5 3 0 0
1965-p2 73 48 47 47 46 12 0 0
1983-p6 53 25 32 28 29 18 0 0
1984-p6 64 32 33 25 24 13 0 0
1985-p6 427 116 116 89 89 84 0 0
1992-p1 91 48 54 35 34 25 0 0
1997-p5 122 51 49 48 51 30 0 0
2022-p2 61 34 30 24 25 25 0 0
2022-p5 265 89 76 80 77 60 0 0
2023-p4 137 52 41 43 45 37 0 0
Total (abs) 1,329 523 504 436 439 316 1 0
Total (%) 100% 39.4% 37.9% 32.8% 33.0% 23.8% 0.1% 0%
Appendix F Distribution of Proof Lengths in LLM-based provers
Figure F1 shows the distribution of the lengths of correct proofs generated by each LLM for the lemmas dataset. We highlight that all LLMs, except o3-mini, struggle to consistently generate lengthy proofs. DeepSeek-Prover-V1.5-RL and Goedel-Prover produce proofs reliably up to a length of 6, while longer proofs become increasingly rare. ReProver, both with and without retrieval, reliably generates proofs of length 3 or shorter. These figures illustrate that the current state-of-the-art models for theorem proving tend to focus on easier and shorter proofs, often overlooking longer proof chains that require planning and a structured understanding of mathematical statements. We note that distribution of human written proofs reflect a high number of problems that require proof chains that go beyond any of the evaluated LLMs. Training a model capable of solving lengthier and more complex proofs remains an ongoing challenge.
(a) DeepSeek-Prover-V1.5-RL
(b) Goedel-Prover-SFT
(c) ReProver (without retrieval)
(d) ReProver (with retrieval)
(e) o3-mini
(f) Lemmas Dataset
Figure F1: Distribution of proof lengths generated by various LLM-based provers on the lemmas dataset (a-e). Histogram (f) illustrates the distribution of length of human-written proofs.
Appendix G Redundancy in LLM generated proofs
In Section 3.3, we refer to proofs that contain unnecessary or redundant structures as “bloated.” Figure G2 illustrates two cases of bloated proofs produced by LLMs. To better reflect the capabilities of modern theorem provers, we manually corrected such proofs when evaluating the length of generated proofs. This correction was done, merely, by commenting out the unnecessary lines, not by revising the proof sketch or by replacing existing with more compact lines.
In Figure G2(a), the proof can be concluded in a single line using omega; however, the LLM introduces an unnecessary structure before reaching this conclusion. In Figure G2(b), the LLM opts to prove hypotheses that are already provided as conditions in the formal statement of the problem. Instead of using them directly to close the goal, the LLM unnecessarily proves them again and then uses the newly derived hypotheses to conclude the proof. We count this as an unnecessary structure within the generated proof. The corrected proofs are presented to the right of the corresponding “bloated” proofs.
(a) Example #1
(b) Example #2
Figure G2: Diagrams that illustrate manual redundancy correction in LLM-generated proofs. The overall proof structure remains intact, with only unnecessary elements/structures removed.
Appendix H Datasheet
Following the framework in Gebru et al. (2021) and similar to Paster et al. (2023), Table LABEL:tbl:datasheet provides the data sheet for our dataset.
Table H5: Datasheet for our dataset
@@ Appendix H Datasheet
How many instances are there in total? 1,329 lemmas.
@@
How was the data associated with each instance acquired? 13 IMO problems were chosen as described in Table 1. The formal proof for each of these problems was written manually by the authors. These proofs were broken into their building blocks leading to a set of 1,329 lemmas written in Lean 4. The proofs for these lemmas were written by the authors. Some of the lemmas that can be proved by automated solvers in Lean were excluded from the dataset. The said automatic solvers are the following: hint, linarith, exact?, simp, omega, ring, norm_cast and norm_num.
What mechanisms or procedures were used to collect the data? Dataset was created by the authors based on the mathematical proofs of 13 IMO problems listed in Table 1.
@@
Was any preprocessing/cleaning/labeling of the data done? No. There was a post processing which excluded some of the lemmas that could be proved directly via automatic solvers in Lean. These lemmas were deemed easy to prove, and since existing automatic solvers can prove them, we do not consider them an obstacle in our larger goal of proving IMO problems.
@@
Is the software that was used to preprocess/clean/label the data available? Lean prover was used to verify the correctness of the process which is an open source system widely used in the field.
@@
Has the dataset been used for any tasks already? We have evaluated the ability of Goedel-Prover, DeepSeek-Prover-V1.5-RL, OpenAI’s o3-mini, Reprover with and without retrieval, Qwen Math 2.5, and Llama 3.1 in proving the lemmas in our dataset.
Is there a repository that links to any or all papers or systems that use the dataset? There will be a link on GitHub which will be public.
@@
Will the dataset be updated? Possibly. If a new version of Lean is released, e.g., Lean 5, we might need to update our dataset based on the new release of Lean and Mathlib. Occasionally, some lemmas in the mathlib library get modified/merged/renamed. Lean automatically marks such items in a proof. We will routinely go over the dataset and fix those issues whenever a new version of Lean is released. The current version of dataset is fully verified in Lean 4.17 which is the latest version.
@@
Appendix I License and authors’ responsibility statement
Appendix J Other motivations
As the Mathlib library grows, and as more lemmas become available in the library, lemmas that can be utilized in writing new proofs, the amount of time for writing a proof for simple problems is expected to reduce. In writing these proofs, mathematicians use their knowledge of Lean and Mathlib to write the proofs. Machines may also be helpful in reducing the time for the formalization of proofs (Buzzard, 2024). In Lean, there are automatic methods that can suggest tactics or even prove some theorems by applying existing lemmas. Researchers also developed LLMs that can suggest tactics aiming to help mathematicians at writing the formal proofs, e.g., the Copilot by Song et al. (2024). However, such LLMs appear to be in their infancy, e.g., Copilot only proves easy lemmas that are designed for new learners of Lean.
On the other hand, for mathematicians, there are always harder theorems and conjectures that can be taken up to be formalized. In other words, there is no limit on how large a mathematical library can grow. Most recently, Kevin Buzzard started a project to formalize the proof for Fermat’s last theorem in Lean, and this project is expected to run for a few years (Buzzard & Taylor, 2024).